OpenStudy (anonymous):

What is the value of s in the equation 3r=10+5s, when r=10 4 8 100 200

Mehek (mehek14):

plug 10 for r \(3*10=10+5s\) \(30=10+5s\) subtract 10 \(20=5s\) divide \(\dfrac{20}{5}\) and you will get the value of s

OpenStudy (anonymous):

Thanks